# A Unified Framework for Input Feature Attribution Analysis

### v1

@@ 6 Conclusion
We introduced a unified evaluation framework for input feature attribution analysis to guide the principled selection of the most suitable explainability technique in practical applications. Our analysis outlines the diverse strengths and trade-offs among TokenEx, TokenIntEx, and SpanIntEx. Our findings particularly underscore SpanIntEx’s superior agreement with human inductive biases and its enhancement of Simulatability, despite its lower comprehensiveness in capturing the model’s reasoning compared to TokenEx and TokenIntEx. This study emphasizes the importance of selecting explanation techniques tailored to specific datasets and model architectures. Future efforts should focus on developing combined methods that enhance both the faithfulness and user comprehensibility of explanations.
@@
Our work introduces a unified framework to evaluate input feature explanations across four key properties. We generated three types of explanations using three attribution methods on two transformer models (BERT-base and BART-base) for two NLU tasks (NLI and fact-checking). This enabled us to assess and compare the properties of each explanation type. Due to computational resource limitations, we did not include larger decoder-only models in our evaluation. Future research could explore these models to provide additional insights. Additionally, examining the properties of different explanations in various tasks beyond NLI and fact-checking would be valuable.
@@
Our findings indicate that span interactive explanations (SpanIntEx) have a notable advantage over other types in Agreement with Human Annotation, Simulatability, and Complexity, suggesting they are easier for humans to understand. This insight could inspire future work to leverage SpanIntEx as the input feature explanation in HCI models. However, SpanIntEx shows low comprehensiveness in faithfulness evaluations. The Louvain algorithm, used for SpanIntEx generation, may limit its comprehensiveness despite using different attribution methods for TokenIntEx. Future work should explore better methods for capturing span interactions and possibly combine SpanIntEx and TokenEx for higher faithfulness, as TokenEx demonstrates a stable advantage in comprehensiveness.
@@ Acknowledgements
Appendix A Detailed Explanation Insertion Method
@@
Appendix B Agreement Dataset Details
To assess how different types of input feature explanations overlap with human annotations, we collected golden explanations of various types from e-SNLI and SpanEx for instances within the SNLI and FEVER datasets, respectively. Detailed information about the annotated explanation types and the number of instances with labeled explanations for these datasets is shown in Table 1. For the SNLI dataset, e-SNLI provides TokenEx explanations, while SpanEx-SNLI includes SpanIntEx explanations and token-level interactions (TokenIntEx explanations). We selected 3,865 overlapping instances and evaluated the human agreement score for different types of input feature explanations. For the FEVER dataset, SpanEx-FEVER includes SpanIntEx and token-level interactions (TokenIntEx explanations). Since no TokenEx explanations are provided, we extracted tokens from the golden TokenIntEx explanations in SpanEx-FEVER as an approximation. These selected instances are also used when evaluating other properties of input feature explanations.
@@ Appendix B Agreement Dataset Details
Table 1: Overview of datasets SNLI (Bowman et al., 2015), FEVER (Thorne et al., 2018), SpanEx (Ray Choudhury et al., 2023) and e-SNLI (Camburu et al., 2018). SpanEx contains instances from SNLI and FEVER datasets, annotated with SpanIntEx explanations including token-level explanations (TokenIntEx explanations). e-SNLI contains instances from SNLI dataset, annotated with TokenEx explanations.
Appendix C Base Model Performance.
As shown in Table 2, we report the performance of fine-tuned BERT-base and BART-base models on the SNLI and FEVER datasets, respectively. These models, fine-tuned for their specific tasks, are used to generate various input feature explanations through different explainability techniques. Importantly, these are the original models that the agent models, as described in §2.4, learn to simulate.
@@
Table 2: The performance of our BERT-base and BART-base models fine-tuned on SNLI and FEVER datasets, respectively, regarding F1 score(%).
Appendix D Explainability Techniques
@@
Shapley. To assign importance scores to each token within the input, we employ the SHAP method following Lundberg and Lee (2017), utilizing Kernel SHAP to approximate Shapley values. For token interactions, we apply Bivariate Shapley  (Masoomi et al., 2023) to assess a token’s influence to model prediction in the presence of another token, particularly when they originate from different parts of the input. The importance score for each token interaction is derived by averaging these mutual influence scores between the involved tokens. With these importance scores for all possible token interactions, we apply the Louvain algorithm  (Ray Choudhury et al., 2023) to identify clusters of neighboring token interactions characterized by dense intra-cluster and sparse inter-cluster relationships, which collectively form span interactions. The importance score for each span interaction is calculated by averaging the importance scores of the token interactions it comprises.
@@
Appendix E Detailed Experiment Results
E.1 Faithfulness
@@
E.2 Agreement with Human Annotation

### v2

@@ 6 Conclusion
We introduced a unified evaluation framework for input feature attribution analysis to guide the principled selection of the most suitable explainability technique in practical applications. Our analysis outlines the diverse strengths and trade-offs among TokenEx, TokenIntEx, and SpanIntEx. Our findings particularly underscore SpanIntEx’s superior performance in Sufficiency, agreement with human inductive biases, its enhancement of Simulatability, and Complexity, compared to TokenEx and TokenIntEx. Future efforts should focus on developing combined methods that enhance all explanation properties.
@@
Our work introduces a unified framework to evaluate input feature explanations across four key properties. We generated three types of explanations using three attribution methods on two Transformer models (BERT-base and BART-base) for two NLU tasks (NLI and fact-checking). Thereby, we can assess and compare the properties of each explanation type. Due to computational resource limitations, we did not include larger decoder-only models in our evaluation. Future research could explore other models to provide additional insights.
We note that our work considered the FEVER and SNLI datasets as they are the only available datasets with annotations of human interactive explanations, required for the Agreement with Human Annotations property. In future work, given the availability of other datasets, examining the properties of different explanations in various tasks beyond NLI and fact checking would be valuable, especially for simpler tasks that consist of only one input part or more complex tasks that consist of more than two parts with possible relationships between them. Additionally, tasks with longer textual inputs, which are known to pose greater challenges for current explainability techniques (Atanasova et al., 2020a), could also be analyzed.
Furthermore, while we consider four widely used explanation properties in this automatic evaluation framework, future works should consider verifying, potentially with supplementary human studies, that the properties are well aligned with the downstream utility of the explanations in different application tasks (Miller, 2019). We note that manual evaluation, while valuable, is time-consuming and costly. Automated evaluation, with our proposed framework, allows for quicker insights, helping prioritize methods that may benefit from human-centered evaluation. Additionally, the properties we evaluate demonstrate why human evaluation is not necessarily required at this stage. Faithfulness measures whether explanations reflect the model’s internal reasoning, a task humans cannot assess (see Faithfulness evaluation guidelines in  Jacovi and Goldberg (2020)). Explanations that fail this test should not be considered for further human evaluation as they can be harmful, e.g. by hiding a model’s flaws and biases. Agreement with Human Annotation already captures alignment with human reasoning, ensuring explanations are plausible. Automated Simulatability correlates strongly with human studies, providing a reliable proxy for replicating the model’s behaviour without the need for expensive human experiments (Pruthi et al., 2022). Finally, automated methods ensure consistency and objectivity, while human annotations can introduce subjectivity and variability. Studies have even shown conflicting results from human evaluations (Poursabzi-Sangdeh et al., 2021; Ribeiro et al., 2016). Automated evaluation provides an objective, reproducible baseline, which can later be supplemented by human evaluations where needed.
We have also employed three base representative explainability methods for each of the three types of input feature explanations. However, more existing and newly emerged base explainability methods could be explored in future work. Additionally, our study focuses solely on post-hoc explainability techniques, while other supervised feature extraction methods could also be investigated (Yu et al., 2021; Liu et al., 2024a). These methods typically treat human-annotated important fragments within the input as gold causal features, akin to our Agreement with Human Annotation measurement. The broader set of properties introduced in our framework could be leveraged to evaluate such explanations more comprehensively. Apart from that, future work could adapt and extend our framework to other forms of explanations such as free-text explanations produced by self-rationalization models  (Liu et al., 2022, 2023b, 2023a, 2024c, 2024b). All said potential future studies are well facilitated by the efficient automated evaluation proposed with our framework.
@@
Our findings indicate that span interactive explanations (SpanIntEx) have a notable advantage over other explanation types in terms of Agreement with Human Annotation, Simulatability, and Complexity, suggesting they are easier for humans to understand. This insight could inspire future work to leverage SpanIntEx as the input feature explanation in HCI models. However, SpanIntEx shows low comprehensiveness in faithfulness evaluations. The Louvain algorithm, used for SpanIntEx generation, may limit its comprehensiveness despite using different attribution methods for TokenIntEx. Future work should explore better methods for capturing span interactions and possibly combine SpanIntEx and TokenEx for higher faithfulness, as TokenEx demonstrates a stable advantage in comprehensiveness.
@@
To ensure a fair comparison, our unified evaluation framework currently considers only the token count differences among various input feature explanations, with interactive explanations flattened. Future work could involve a human-in-the-loop approach to account for the effects of interactive explanations beyond just token count differences. For example, a display system could visually present highlighted tokens and interactions to gauge human preferences. Our work provides a starting point for comparing input feature explanations from an automated evaluation perspective, and future research could explore additional factors, such as psychological elements and visual aspects, from a human perspective, which would benefit more non-expert users.
Another limitation of this work is that we focus solely on the automatic evaluation of input feature explanations without examining the potential biases they may exhibit. For instance, these explanations might favor certain words or phrases from the input sequence, for example, sometimes emphasizing some prepositions that might bear less meaning, raising questions about whether such biases stem from the models themselves or from the explainability techniques used. Also, it is worth checking, especially in sensitive domains such as healthcare and law, how contradictory explanations for different model decisions differed, which might sway the decision-making.
@@ Acknowledgements
Appendix A Unified Faithfulness Evaluation: Explanation Masking Details
As discussed in Section §2.2, we introduce the dynamic threshold $\theta_{x,k}$ to ensure an identical number of tokens from different types of top input feature explanations for the same instance in Unified Faithfulness evaluation.
For the Unified Comprehensiveness evaluation, we conduct a similar process for all three explanation types separately. Using TokenEx as an example, we first calculate the maximum number of top TokenEx for disturbing each instance $x$ as $k_{\texttt{TokenEx}}(x)$ following Eq. 7. To omit the TokenEx tokens from the original input, we replace them with [MASK] tokens, while keeping the rest unchanged. The specific [MASK] token used depends on the model architecture. We then gradually increase the number of TokenEx tokens masked out until it reaches $k_{\texttt{TokenEx}}(x)$ and record the corresponding changes in model predictions. The average prediction change across the dynamic threshold and all instances gives the Unified Comprehensiveness score for TokenEx. For TokenIntEx and SpanIntEx, the only difference is that we mask out token pairs for TokenIntEx and span pairs for SpanIntEx, with the maximum explanations masked out as $k_{\texttt{TokenIntEx}}(x)$ and $k_{\texttt{SpanIntEx}}(x)$ for each instance $x$, calculated using Eq. 6 and Eq. 8 respectively.
For the Unified Sufficiency evaluation, we conduct experiments for the three explanation types separately. Unlike Unified Comprehensiveness, we retain only the tokens/token pairs/span pairs for the input while masking out all other tokens with [MASK] tokens for each instance, depending on the model architecture used. We first calculate the maximum number of top explanations involved in disturbance for each explanation type for instance $x$ using Eq. 6, Eq. 7, and Eq. 8. Then, we keep the token/token pairs/span pairs in the model input by masking out all other tokens, starting with one explanation and adding one more explanation for each subsequent disturbance until the total number of explanations reaches $k_{\texttt{TokenEx}}(x)$, $k_{\texttt{TokenIntEx}}(x)$, or $k_{\texttt{SpanIntEx}}(x)$. Meanwhile, we record the model predictions for each disturbance. The Unified Sufficiency score for each explanation type is then calculated by averaging the prediction changes across the dynamic threshold for that explanation type, considering all instances.
Appendix B Detailed Explanation Insertion Method
@@
Appendix C Agreement Dataset Details
To assess how different types of input feature explanations overlap with human annotations, we collected golden explanations of various types from e-SNLI and SpanEx for instances within SNLI and FEVER, respectively. Detailed information about the annotated explanation types and the number of instances with labeled explanations for these datasets is shown in Table 1. For the SNLI dataset, e-SNLI provides TokenEx explanations, while SpanEx-SNLI includes SpanIntEx explanations and token-level interactions (TokenIntEx explanations). We selected 3,865 overlapping instances and evaluated the human agreement score for different types of input feature explanations. For FEVER, SpanEx-FEVER includes SpanIntEx and token-level interactions (TokenIntEx explanations). Since no TokenEx explanations are provided, we extracted tokens from the golden TokenIntEx explanations in SpanEx-FEVER as an approximation. These selected instances are also used when evaluating other properties of input feature explanations.
@@ Appendix C Agreement Dataset Details
Table 1: Overview of datasets SNLI (Bowman et al., 2015), FEVER (Thorne et al., 2018), SpanEx (Ray Choudhury et al., 2023) and e-SNLI (Camburu et al., 2018). SpanEx contains instances from SNLI and FEVER, annotated with SpanIntEx explanations including token-level explanations (TokenIntEx explanations). e-SNLI contains instances from SNLI dataset, annotated with TokenEx explanations.
Appendix D Base Model Performance.
As shown in Table 2, we report the performance of fine-tuned BERT-base and BART-base models on SNLI and FEVER, respectively. These models, fine-tuned for their specific tasks, are used to generate various input feature explanations through different explainability techniques. Importantly, these are the original models that the agent models, as described in §2.4, learn to simulate.
@@
Table 2: The performance of our BERT-base and BART-base models fine-tuned on SNLI and FEVER, respectively, regarding F1 score(%).
Appendix E Explainability Techniques
@@
As denoted in Section §2.1, $x_{i}$ represents the $i$th token with instance $x$. To better illustrate the explainability techniques below, we use $F$ as the set of all tokens within this instance and $S$ as the subset of $F$. All explanations are obtained using model $M$, which is omitted in the following notions for brevity. We use $A_{\texttt{TokenEx}}(x_{i})$ to denote the attribution score generated by explainability technique $A$ for the $i$th token $x_{i}$, $A_{\texttt{TokenIntEx}}(x_{i},x_{j})$ as the attribution score for token interaction $(x_{i},x_{j})$, $A_{\texttt{TokenIntEx}}(x_{i}\mid x_{j})$ as the importance score of token $x_{i}$ conditioned on $x_{j}$ is present when the directed importance between tokens within $(x_{i},x_{j})$ is considered in some attribution techniques, $A_{\texttt{SpanIntEx}}(span_{i}^{0},span_{i}^{1})$ as the attribution score for corresponding span interaction, where $span_{i}^{0}=(x_{s},...,x_{s+l_{1}})$ is a span from part1 and $span_{i}^{1}=(x_{t},...,x_{t+l_{2}})$ is a span from part2 of the input. Note that in Section §2.1, we use $a^{x}_{k}$ to denote the importance score of the $k$th most important explanation of instance $x$; here, we only focus on the attribution scores of explanations without ranking them.
Shapley. For TokenEx, we employ the SHAP method to assign importance scores to each token within the input by removing each token separately and computing its removal effect on the model prediction with different subsets of other tokens presented to the model, following Lundberg and Lee (2017).
$\begin{split}Shap_{\texttt{TokenEx}}(x_{i})&=\sum_{S\subseteq F\setminus\{x_{i}\}}\frac{|S|!(|F|-|S|-1)!}{|F|!}\\ &\quad[f(S\cup\{x_{i}\})-f(S)]\end{split}$ (21)
$Shap_{\texttt{TokenEx}}(x_{i})$ denotes the importance score of token $x_{i}$. As the calculation of $Shap_{\texttt{TokenEx}}(x_{i})$ is computationally expensive, we utilize Kernel SHAP to approximate these Shapley values.
For TokenIntEx, we first apply Bivariate Shapley  (Masoomi et al., 2022) to assess the mutual importance scores between two tokens, which are from different parts of the input, within a token interaction, and then average these two mutual importance scores as importance score of this token interaction. Specifically, to compute the importance score of a token $x_{i}$ conditioned on the presence of token $x_{j}$, the sets of tokens $S$ considered are limited to those containing token $x_{j}$, while the impact of other sets of tokens influences the importance of $x_{i}$ is ignored in this case. Thus, $Shap_{\texttt{TokenIntEx}}(x_{i}\mid x_{j})$ can be calculated by:
$\begin{split}Shap_{\texttt{TokenIntEx}}(x_{i}\mid x_{j})&=\sum_{x_{j}\in S\subseteq F\setminus\{x_{i}\}}\frac{|S|!(|F|-|S|-1)!}{|F|!}\\ &\quad[f(S\cup\{x_{i}\})-f(S)]\end{split}$ (22)
The importance score $Shap_{\texttt{TokenIntEx}}(x_{i},x_{j})$ for token interaction $(x_{i},x_{j})$ are calculated by averaging $Shap_{\texttt{TokenIntEx}}(x_{i}\mid x_{j})$ and $Shap_{\texttt{TokenIntEx}}(x_{j}\mid x_{i})$. We also use Kernel Shapley to approximate the calculation of Bivariate Shapley value.
For SpanIntEx, we first apply the Louvain Community Detection algorithm  (Blondel et al., 2008) to extract span interactions and then average the importance scores of token interactions comprised in each span interaction as its importance score, following Ray Choudhury et al. (2023).
To extract span interactions, we first construct a directed bipartite graph for instance $x$, by taking each token $x_{i}$ from the input as node $i$ and the mutual importance scores between each two tokens from different parts obtained above as the weights of directed edges connecting them. Louvain Community Detection algorithm is then applied to search for communities of nodes with dense intra-cluster and sparse inter-cluster relationships. With each community of nodes(tokens) $S_{p}$ obtained, we can get one span interaction $(span_{p}^{0},span_{p}^{1})$, where the two spans consist of neighboring tokens from the part1 subset of this community $S_{p}^{0}$, and the part2 subset of it $S_{p}^{1}$ respectively.
Then we calculate the importance score of this span interaction by averaging the importance scores of all token interactions it comprises.
$Shap_{\texttt{SpanIntEx}}(span_{p}^{0},span_{p}^{1})$ $=$ (23)
$\quad\frac{\sum\limits_{i,j}^{\begin{array}[]{l}\begin{subarray}{c}x_{i}\in S_{p}^{0}\\ x_{j}\in S_{p}^{1}\end{subarray}\end{array}}Shap_{\texttt{TokenIntEx}}(x_{i},x_{j})}{|S_{p}^{0}\cup S_{p}^{1}|}$
Note that in the following, no matter which explainability techniques to assign importance score to TokenIntEx, we apply the same method as stated above to extract span interactions and compute their importance scores, $A_{\texttt{SpanIntEx}}(span_{p}^{0},span_{p}^{1})$, based on corresponding token interaction importance score, $A_{\texttt{TokenIntEx}}(x_{i},x_{j})$.
@@
Appendix F Detailed Experiment Results
F.1 Faithfulness
@@
F.2 Agreement with Human Annotation